# GDSG: Graph Diffusion-based Solution Generation for Optimization Problems in MEC Networks

### v1

@@ Abstract
Optimization is crucial for MEC networks to function efficiently and reliably, most of which are NP-hard and lack efficient approximation algorithms. This leads to a paucity of optimal (ground-truth) data, constraining the effectiveness of conventional deep learning approaches. Most existing learning-based methods necessitate extensive optimal data and fail to exploit the potential benefits of suboptimal data that can be obtained with greater efficiency and effectiveness. Taking the multi-server multi-user computation offloading (MSCO) problem, which has been widely observed in systems such as Internet-of-Vehicles (IoV) and Unmanned Aerial Vehicle (UAV) networks, as a concrete scenario, we present a Graph Diffusion-based Solution Generation (GDSG) method. This approach is designed to work with suboptimal datasets while converging to the optimal solution large probably. We transform the network optimization issue into a distribution-learning problem and offer a clear explanation of learning from suboptimal training datasets. We build GDSG as a multi-task diffusion generative model utilizing a Graph Neural Network (GNN) to acquire the distribution of high-quality solutions. We use a simple and efficient heuristic approach to obtain a sufficient amount of training data composed entirely of suboptimal solutions. In our implementation, we enhance the backbone GNN and achieve improved generalization. GDSG also reaches nearly 100% task orthogonality, ensuring no interference between the discrete and continuous solution generation training objectives. We further reveal that this orthogonality arises from the diffusion-related training loss in GDSG, rather than the neural network architecture itself. The experiments demonstrate that GDSG surpasses other benchmark methods on both the optimal and suboptimal training datasets. Concerning the minimization of computation offloading costs, GDSG can achieve savings of up to $56.63\%$ on the ground-truth training set and $41.07\%$ on the suboptimal training set compared to the existing discriminative methods. The MSCO datasets has open-sourced at http://ieee-dataport.org/13824, as well as the GDSG algorithm codes at https://github.com/qiyu3816/GDSG.
@@
Mobile edge computing, network optimization, computation offloading, generative AI, graph diffusion.
@@ I Introduction
In various cutting-edge scenarios of Mobile Edge Computing (MEC) [1, 2], optimization demands are ubiquitous, such as trajectory optimization in Unmanned Aerial Vehicle (UAV) networks [3, 4, 5, 6], utility and latency optimization in edge networks for Federated Learning (FL) [7, 8], sum-rate and coverage optimization in Reconfigurable Intelligent Surface (RIS)-enhanced networks [9, 10, 11, 12, 13], efficiency optimization in Internet-of-Vehicles (IoV) [14, 15, 16, 17], and other resource allocation tasks [18, 19, 20, 21, 22, 23, 24, 25]. Despite the remarkable achievements of both numerical algorithms and Artificial Intelligence (AI) methods, a key challenge remains that a significant portion of these problems are NP-hard, with no efficient approximation algorithms currently available. This not only limits algorithms to suboptimal results, despite intensive manual design efforts, but also curtails the potential of AI-based methods. AI approaches typically require large-scale, optimally solved examples, which is unrealistic for problems that are inherently difficult to solve.
Fortunately, the success of generative AI has demonstrated its ability to generalize beyond the training data, offering a more robust AI approach to data quality. Large Language Models (LLMs) trained on limited texts have been shown to generate novel outcomes in response to diverse queries [26, 27], rather than merely reassembling the training data. Similarly, Generative Diffusion Models (GDMs) trained on image datasets can produce significantly different images under varying conditions [28, 29], rather than simply recalling training samples. Based on this, can we reasonably conclude that AI models for optimization do not necessarily need to rely on data composed of optimal solutions?
For many optimization problems in MEC that lack efficient solution methods, suboptimal solutions can often be more easily derived for many of these problems, while obtaining a sufficient number of optimal solutions is time-consuming and labor-intensive. For example, the multi-objective nature of some optimization problems [3, 7, 12, 15, 22, 19, 20, 23, 21, 11, 18, 10, 16, 17, 5, 6, 4] can be tackled through divide-and-conquer strategies to obtain suboptimal solutions, which are not only simple to implement but also computationally efficient. The core idea is to piece together existing simple algorithms to construct a complete suboptimal solution, which can obtain references from numerous previous studies. Many of these methods also feature low complexity, making them suitable for constructing large-scale datasets. While these suboptimal instances may not be very beneficial for current AI approaches, our work will leverage GDMs from generative AI to explore information and sample better solutions, with the potential to generate optimal solutions.
@@ I-A Motivation and Contribution
To address the aforementioned challenges, we attempt to use generative AI to learn from suboptimal datasets for solution generation in MEC network optimization problems. This is an approach that has not been explored before, and we will provide a detailed analysis of its effectiveness, along with experimental results.
We introduce a framework designed to acquire a parameterized solution distribution from imperfect solution samples. This is similar to the high-quality solution distribution in the DIMES [30], but it uses Deep Reinforcement Learning (DRL) instead of generative AI for optimization. Essentially, we view the feasible solution space of a given input as a distribution of solutions, with the optimal solution having the highest likelihood within this distribution. Subsequently, we parameterize the resulting solution distribution to facilitate distribution learning. Following this step, the parameterized distribution derived from the input can be concurrently sampled along with the approximation to the optimal solution. Distribution learning is an inherent capability of generative AI, with GDMs being among the most efficient [31], supporting various types of data patterns. To introduce a novel GDM-based approach for generating solutions to optimization problems and support as many applicable scenarios as possible, we consider the close relationship between optimization problems in MEC networks and graph-structured data, such as topology and routing trajectories [19, 18, 25, 20, 12]. Accordingly, we propose Graph Diffusion-based Solution Generation (GDSG), a method that solves problems at the graph level. Taking into account the multi-server multi-user setting in MEC networks, which may arise in scenarios like UAV and IoV networks, we model a multi-server multi-user computation offloading problem. We construct a suboptimal training set using heuristic methods and validate the performance of GDSG on this problem.
@@
To the best of our knowledge, this is the first research attempt to build GDM-enabled solution generation method considering the training data limitations of complex optimization problems in MEC networks. The key contributions of our work are outlined below.
  • We delve into the potential of suboptimal data samples and elucidate the rationale behind learning the solution distribution. We provide discussion about the convergence to the optimal solution and related conditions, with some insightful references summarized.
  • We propose a graph optimization approach, trained solely on suboptimal data samples, demonstrating the capability to generate high-quality solutions and achieve optimal solutions for some inputs through parallel sampling. This alleviates the rely on laborious preparation of an optimal solution training set.
@@
  • We provide some practical insights from engineering implementation. By explicitly improving the Graph Neural Network (GNN) gating mechanism, the model’s generalization ability across different graph scales is enhanced. Additionally, our analysis of multi-task simultaneous training reveals the orthogonality introduced by the diffusion model’s loss objective, which results in negligible interference between tasks.
@@ I-B Related works
I-B1 Learning-based optimization methods
Current learning-based approaches to network optimization can be categorized based on their reliance on training data. The first category utilizes ground-truth datasets [22, 23, 21, 25, 20, 5, 9], either created from scratch using standard solvers, brute force techniques, or directly drawn from open-source datasets. Nevertheless, open-source datasets have limited coverage, and the solvers’ reliance on problem-specific assumptions and the high cost of dataset construction renders it impractical for new scenarios. The second category aims to reduce reliance on already solved datasets [17, 18, 19, 7, 6, 15, 14]. This approach primarily relies on reinforcement learning, guiding the model to explore improved policies through carefully crafted rewards and losses. However, this method heavily depends on expert knowledge and necessitates adjustments when the scenario changes.
@@ I-B2 Graph diffusion and optimization
As a common type of generative model, diffusion models involve a process of introducing noise and then learning to eliminate it. This method enables the model to predict and gradually recover the desired distribution by cyclically rectifying noise. An illustrative instance is the Denoising Diffusion Probabilistic Model (DDPM) [32], capable of producing distributions without constraints, while its variations can generate distributions based on specific conditions [33, 34].
Graph diffusion, on the other hand, involves learning and generating non-Euclidean graph-structured data by utilizing GNNs [35]. This method has shown success in solving combinatorial optimization problems like the Traveling Salesman Problem (TSP) and the Maximum Independent Set problem (MIS) with impressive outcomes [36, 37]. Specifically, T2T [37] enhances the sampled solutions by integrating gradient-guided generation based on DIFUSCO [36]. To date, graph diffusion models have proven beneficial by explicitly modeling graph structures to learn multi-modal distribution features.
Recent study [38, 39, 40] has delved into integrating diffusion models with network optimization, albeit primarily through practical cases in various optimization problems without concern for data robustness. Some works [41, 42, 43] also leverage GDMs as a stronger policy network in DRL, an emerging approach in the AI community, to address network optimization problems.
Different from the related works above, the proposed GDSG approach provides the implementation of graph-level modeling in complex optimization solving for MEC networks. Especially, our research leads the way in investigating the potential of suboptimal datasets utilized by GDMs and achieves a robust performance. Creating suboptimal datasets can be efficiently accomplished using basic heuristic algorithms which ensures ensures the training data for GDSG. Based on the generative model’s understanding of the solution distribution, optimal solution generation can be attainable through parallel sampling. Out implementation also allows multi-task optimization and enhanced generalization.
@@ II System Model and Problem Formulation
With the development of systems such as UAV networks, RIS deployments, and IoV networks, the demand for more complex computational tasks has led to resource shortages on user devices. Computation offloading [2] has become an essential solution, and with the increasing number of edge devices, multi-server and multi-user offloading scenarios are becoming increasingly common [24, 5, 8, 7, 10, 6].
@@ II-A System Model
Our system consists of edge servers and mobile users, where mobile users represent a unified abstraction of devices such as smartphones, drones, and vehicles. The number of servers and users in the system is variable, allowing for dynamic node mobility and the flexible joining or leaving of nodes based on task requirements.
II-A1 System components
@@ II-A3 Binary offloading
During the offloading optimization, each user is assumed to have an indivisible computing task, either executed locally or offloaded to a MEC server for execution. To intuitively formalize the relationship between edges and nodes, we define a binary function
@@
$\forall i\in\{1,...,L\},\ \forall j\in\{1,...,M\}$
to determine whether $U_{j}$ is the source node of the directed edge $E_{i}$. Therefore, the outgoing edges of a user node $U_{j}$ should satisfy $\sum_{\Upsilon(E_{i},U_{j})=1}D_{i}\leq 1$.
In addition, the total amount of computational resources allocated to the tasks offloaded to a server cannot exceed the available computational resources of the server. We define a function
@@
$\forall i\in\{1,...,L\},\ \forall k\in\{1,...,K\}$
to determine whether $S_{k}$ is the destination node of the directed edge $E_{i}$. Therefore, the incoming edges of a server node $S_{k}$ should satisfy $\sum_{\Omega(E_{i},S_{k})=1}D_{i}A_{i}\leq 1$.
@@ II-B Weighted Cost of Delay and Energy Consumption
II-B1 Parameters of a user node for local computation
Let $\mathbf{I}=\{I_{1},...,I_{L}\}$, $\mathbf{V}=\{V_{1},...,V_{L}\}$ and $\mathbf{F}=\{F^{l}_{1},...,F^{l}_{L}\}$ be the input data sizes [bits], required computational resources [cycles] and available local computational resources [cycles per second] of the user tasks on each edge. Let $\mathcal{\alpha}=\{\alpha_{1},...,\alpha_{L}\}$ represent the weight of the task cost, where $\alpha_{i}\in[0,1]$. So we have $\alpha\to 1$ if the delay cost is emphasized while the energy consumption cost is emphasized if $\alpha\to 0$. Clearly, the parameters related to local computation are the same for all outgoing edges of the same user node. So we have
@@ II-B1 Parameters of a user node for local computation
II-B2 Cost of local execution
@@
$\mathcal{C}^{l}_{i}=\alpha_{i}\tau^{l}_{i}+(1-\alpha_{i})\omega^{l}_{i},\ \forall i\in\{1,...,L\}.$ (4)
II-B3 Cost of transformation for offloading
Let =$\mathcal{H}=\{h_{1},...h_{i},...,h_{L}\},h_{i}\in[0,1]$ be the channel gain and $B$ be the channel bandwidth of the wireless connection represented by each edge. Let the power [W] of the uplink be $P_{t}$ with additive Gaussian white noise of zero mean and variance $N_{0}$. Then the uplink received signal-to-interference-plus-noise ratio (SINR) is $SINR_{i}=\frac{P_{t}(h_{i})^{2}}{N_{0}+P_{t}\sum(h_{j})^{2}}$ where $E_{j}$ and $E_{i}$ has the same server node. In this way, the transmission rate of $E_{i}$ is calculated as $r^{u}_{i}=B\log_{2}(1+SINR_{i})$. The weighted cost of transmission can be calculated as
$\mathcal{C}^{t}_{i}=\alpha_{i}\frac{I_{i}}{r^{u}_{i}}+(1-\alpha_{i})P_{t}\frac{I_{i}}{r^{u}_{i}},\ \forall i\in\{1,...,L\}.$ (5)
II-B4 Cost of offloading execution

### v3

# GDSG: Graph Diffusion-based Solution Generator for Optimization Problems in MEC Networks

@@ Abstract
Optimization is crucial for the efficiency and reliability of multi-access edge computing (MEC) networks. Many optimization problems in this field are NP-hard and do not have effective approximation algorithms. Consequently, there is often a lack of optimal (ground-truth) data, which limits the effectiveness of traditional deep learning approaches. Most existing learning-based methods require a large amount of optimal data and do not leverage the potential advantages of using suboptimal data, which can be obtained more efficiently. To illustrate this point, we focus on the multi-server multi-user computation offloading (MSCO) problem, a common issue in MEC networks that lacks efficient optimal solution methods. In this paper, we introduce the graph diffusion-based solution generator (GDSG), designed to work with suboptimal datasets while still achieving convergence to the optimal solution with high probability. We reformulate the network optimization challenge as a distribution-learning problem and provide a clear explanation of how to learn from suboptimal training datasets. We develop GDSG, a multi-task diffusion generative model that employs a graph neural network (GNN) to capture the distribution of high-quality solutions. Our approach includes a straightforward and efficient heuristic method to generate a sufficient amount of training data composed entirely of suboptimal solutions. In our implementation, we enhance the GNN architecture to achieve better generalization. Moreover, the proposed GDSG can achieve nearly 100% task orthogonality, which helps prevent negative interference between the discrete and continuous solution generation training objectives. We demonstrate that this orthogonality arises from the diffusion-related training loss in GDSG, rather than from the GNN architecture itself. Finally, our experiments show that the proposed GDSG outperforms other benchmark methods on both optimal and suboptimal training datasets. Regarding the minimization of computation offloading costs, GDSG achieves savings of up to 56.62% on the ground-truth training set and 41.06% on the suboptimal training set compared to existing discriminative methods [^1] The MSCO datasets are open-sourced at https://dx.doi.org/10.21227/386p-7h83, and the GDSG algorithm codes at https://github.com/qiyu3816/GDSG..
@@
Multi-access edge computing, network optimization, computation offloading, generative AI, graph diffusion.
@@ I Introduction
In various cutting-edge scenarios of multi-access edge computing (MEC) [1, 2], optimization demands are ubiquitous, such as trajectory optimization in uncrewed aerial vehicle (UAV) networks [3, 4, 5, 6, 7, 8, 9], utility and latency optimization in edge networks for federated learning (FL) [10, 11, 12], sum-rate and coverage optimization in reconfigurable intelligent surface (RIS)-enhanced networks [13, 14, 15, 16, 17], efficiency optimization in Internet-of-Vehicles (IoV) [18, 19, 20], and other resource allocation tasks [21, 22, 23, 24, 25, 26, 27]. Despite the significant advancements in numerical algorithms and artificial intelligence (AI) methods, a key challenge persists: a large number of these problems are NP-hard, and there are currently no efficient approximation algorithms available. This limitation not only restricts algorithms to producing suboptimal results, despite extensive manual design efforts, but also hinders the potential of AI-based methods, which typically rely on large datasets consisting of optimally solved examples that are inherently difficult to obtain in practice.
Fortunately, the success of generative AI has demonstrated its ability to generalize beyond the training data, offering a more robust AI approach to data quality. Large language models (LLMs) trained on a limited amount of text can generate novel responses to diverse queries [28, 29] rather than simply reassembling the training data. Similarly, generative diffusion models (GDMs) trained on image datasets can create significantly different images under various conditions [30, 31], rather than simply recalling training samples. Based on this evidence, we propose the hypothesis that generative AI models can solve complex network optimization problems, even in the absence of an exclusively optimal dataset.
@@ I-A Motivation and Contribution
Many optimization problems in MEC networks lack efficient solution methods. Existing commercial and open-source solvers, such as Gurobi [32], Mosek [33], CPLEX [34], and GEKKO [35], often face excessively high algorithmic complexity during iterative optimization processes. For common challenges, such as non-convex optimization and mixed-integer programming, these solvers can typically guarantee efficiency only in low-dimensional cases, sometimes with fewer than $10$ dimensions. Furthermore, many network optimization problems still lack mature and effective algorithms, leaving them without ready-made solvers. The limitations of traditional solvers in handling high-complexity problems highlight the necessity for AI-based methods. However, generating a large number of optimal solution samples for training purposes is prohibitively expensive, both in terms of time and computational resources. In such scenarios, deriving suboptimal solutions often proves easier than obtaining a sufficient number of optimal solutions, which can be time-consuming and labor-intensive. For example, the multi-objective nature of certain optimization problems can be effectively addressed using divide-and-conquer strategies. This approach enables the development of suboptimal solutions that are straightforward to implement and computationally efficient, making them suitable for constructing large-scale datasets. Although these suboptimal instances may not offer significant advantages for current AI techniques, they provide an opportunity to explore the potential of GDM in generating solution distributions and achieving improved outcomes through parallel sampling.
To validate our earlier hypothesis, we employ generative AI to extract insights from suboptimal datasets to develop solutions for MEC optimization problems. In this context, we consider the feasible solution space of a given input as a distribution of potential solutions, with the optimal solution being the one that has the highest probability within this distribution. We then parameterize the resulting solution distribution to facilitate effective distribution learning. Following this step, we can simultaneously sample from the parameterized distribution generated from the input, along with an approximation of the optimal solution. Generative models possess the ability to learn a target distribution and sample from it. Among these models, GDM stands out for the balance of performance and efficiency [36]. Consequently, we utilize GDM to learn the parameterized solution distribution and to further sample high-quality solutions from it.
Given that many problems in MEC can be formulated as graph optimization problems, such as topology design and routing scheduling [22, 21, 26, 37, 16]. To address this, we focus on the multi-server multi-user computation offloading (MSCO) problem and utilize graph diffusion techniques to find a solution. We propose a solution generator called the graph diffusion-based solution generator (GDSG). To train GDSG, we first create a suboptimal training set using heuristic methods and then employ a conditional generation mechanism. During the solving phase, we use the model to perform parallel sampling of high-quality solutions, which helps us approximate the optimal solution.
@@
To the best of our knowledge, this is a pioneering attempt to solve complex optimization problems that take into account the training data limitations in MEC networks. The key contributions of our work are outlined below.
  • We explore the potential of utilizing suboptimal data samples and clarify the reasoning for learning the solution distribution instead of relying solely on GDM with a large number of optimal solution samples. Our theoretical framework discusses the convergence toward the optimal solution and summarizes the conditions required for convergence.
  • We present GDSG, a method trained on suboptimal data for graph optimization, demonstrating the capability to generate high-quality solutions with specific inputs through parallel sampling. This approach offers a more efficient network optimization solution than numerical algorithms while also incurring lower data costs compared to traditional AI methods, all while ensuring optimal convergence.
@@
  • We provide some practical insights from engineering implementation aspects. By explicitly enhancing the graph neural network (GNN) gating mechanism, we improve the model’s generalization ability across different graph scales. Additionally, our analysis of multi-task training reveals the orthogonality introduced by the diffusion model’s loss objective, leading to negligible interference between tasks.
@@ I-B Related works
I-B1 Learning-based optimization
Current learning-based approaches to network optimization can be classified based on their reliance on training data. The first category utilizes ground-truth datasets [38, 25, 24, 26, 37, 5, 13], either created from scratch using standard solvers, brute force techniques, or derived from open-source datasets. However, open-source datasets often have limited coverage, and the use of solvers based on problem-specific assumptions, along with the high cost of dataset construction, makes this approach impractical for new scenarios. The second category seeks to minimize reliance on already solved datasets [19, 21, 22, 10, 6, 18]. This approach primarily uses reinforcement learning, where the model is guided to discover improved policies through the careful design of rewards and losses. Nevertheless, this method heavily depends on expert knowledge and requires adjustments when the scenarios change.
@@ I-B2 Graph diffusion and optimization
Diffusion models are a common type of generative model that involve introducing noise into data and then learning to remove that noise. This approach enables the model to predict and progressively restore the desired data distribution by repeatedly correcting for the noise. One notable example is the denoising diffusion probabilistic model (DDPM) [39], which can produce distributions based on specific conditions [40, 41].
Graph diffusion, on the other hand, involves learning and generating non-Euclidean graph-structured data using GNNs [42]. This approach has demonstrated success in addressing combinatorial optimization problems such as the traveling salesman problem (TSP) and the maximum independent set problem (MIS), yielding impressive results in both [43] and [44]. Specifically, the T2T method [44] improves the sampled solutions by incorporating gradient-guided generation based on DIFUSCO [43]. Thus far, graph diffusion models have proven advantageous by explicitly modeling graph structures to learn features of multi-modal distributions.
Recent studies [45, 46, 47] have explored the integration of diffusion models with network optimization. However, this research mainly focuses on practical applications in various optimization problems without considering data robustness. Additionally, with new combinations of GDM and reinforcement learning in the AI-community, some studies [48, 49, 50] utilize GDM as a more effective policy network in deep reinforcement learning (DRL) to tackle network optimization challenges. Unlike the related works mentioned above, the proposed GDSG approach provides an effective implementation of graph-level modeling for solving complex problems via multi-task optimization in MEC networks. The suboptimal datasets can be efficiently created using basic heuristic algorithms, which ensure the quality of the training data for GDSG. Our research specifically explores the potential of utilizing suboptimal datasets with GDMs, which can lead to robust performance outcomes. Our theoretical findings and engineering results provide GDSG with interpretable superior performance and reduce implementation barriers. This approach helps to eliminate confusion arising from unknown underlying principles and addresses engineering challenges associated with data scarcity.
@@ II System Model and Problem Formulation
The advancement of various MEC systems [11, 10, 51, 52, 53] has increased the demand for complex computational tasks, resulting in resource shortages on user devices. In this context, computation offloading has emerged as a crucial solution to address this issue.
@@ II-A System Model
Our considered MSCO system consists of multiple edge servers and mobile users, where mobile users represent a unified abstraction of devices such as smartphones, drones, and vehicles. The number of servers and users in the system is variable, allowing for dynamic node mobility and the flexible joining or leaving of nodes based on task requirements.
II-A1 System overview
@@ II-A3 Binary offloading
During the offloading optimization, each user is assumed to have an indivisible computing task, either executed locally or offloaded to a MEC server for execution. To intuitively formalize the relationship between edges and nodes, we define a binary function, i.e.,
@@
$\forall i\in\{1,...,L\},\ \forall j\in\{1,...,M\}$ ,
to determine whether $U_{j}$ is the source node of the directed edge $E_{i}$ or not. Therefore, the outgoing edges of a user node $U_{j}$ should satisfy $\sum_{\Upsilon(E_{i},U_{j})=1}D_{i}\leq 1$.
In addition, the total amount of computational resources allocated to the tasks offloaded to a server cannot exceed the available computational resources of the server. Therefore, we define a function
@@
$\forall i\in\{1,...,L\},\ \forall k\in\{1,...,K\}$ ,
to determine whether $S_{k}$ is the destination node of the directed edge $E_{i}$ or not. As a result, the incoming edges of a server node $S_{k}$ should satisfy $\sum_{\Omega(E_{i},S_{k})=1}D_{i}A_{i}\leq 1$.
@@ II-B Weighted Cost of Delay and Energy Consumption
Let $\mathbf{I}=\{I_{1},...,I_{L}\}$, $\mathbf{V}=\{V_{1},...,V_{L}\}$ and $\mathbf{F}=\{F^{l}_{1},...,F^{l}_{L}\}$ be the input data sizes [bits], required computational resources [cycles] and available local computational resources [cycles per second] of the user tasks on each edge. Let $\mathcal{\alpha}=\{\alpha_{1},...,\alpha_{L}\}$ represent the weight of the task cost, respectively, where $\alpha_{i}\in[0,1]$. So we have $\alpha\to 1$ if the delay cost is emphasized while the energy consumption cost is emphasized if $\alpha\to 0$. Clearly, the parameters related to local computation are the same for all outgoing edges of the same user node. Thus, we have
@@
II-B1 Local execution cost
@@
$C^{l}_{i}=\alpha_{i}\tau^{l}_{i}+(1-\alpha_{i})\omega^{l}_{i},\ \forall i\in\{1,...,L\}.$ (4)
II-B2 Offloading execution cost
Let =$\mathcal{H}=\{h_{1},...,h_{i},...,h_{L}\},\ h_{i}\in[0,1]$ be the channel gain and $B$ be the channel bandwidth of the wireless connection represented by each edge. Let the power [W] of the uplink be $P_{t}$ with additive Gaussian white noise of zero mean and variance $N_{0}$. Then the uplink received signal-to-interference-plus-noise ratio (SINR) is given as $SINR_{i}=\frac{P_{t}(h_{i})^{2}}{N_{0}+P_{t}\sum(h_{j})^{2}}$, where $E_{j}$ and $E_{i}$ has the same server node. In this way, the transmission rate of $E_{i}$ is calculated as $r^{u}_{i}=B\log_{2}(1+SINR_{i})$. The weighted cost of uploading transmission can be calculated as
${C}^{t}_{i}=\alpha_{i}\frac{I_{i}}{r^{u}_{i}}+(1-\alpha_{i})P_{t}\frac{I_{i}}{r^{u}_{i}},\ \forall i\in\{1,...,L\}.$ (5)